if you havw 100 coins totalling 7.00.no nickels.what coins do you have if 60% of the coins are the one of a kind....
Answer by josgarithmetic(39620)

Here is most of what the process can be.
p, d, q for pennies dimes quarters.

p+d+q=100 and p+10d+25q=700. Remember 60 of the 100 coins will be one kind.

Subtract the coin count equation from the money count equation:
9d+24q=600

Either d will be 60 OR q will be 60 but NOT BOTH. Which choice gives you whole numbers for both d and q ?